OpenStudy (trent):

The ratio of men to women working for a company 3 to 5 . If there are 51 men working for the company, what is the total number of employees?

OpenStudy (anonymous):

Men / Women = 3/5 Women = Men * 5 / 3

OpenStudy (tcp2315):

136

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

Sorry. I thought you were giving number of women.

OpenStudy (tcp2315):

hahah its ok. trent to clairfy there are 85 women and 136 in total

OpenStudy (trent):

how did you come up with that how would i write that out (thank you)

OpenStudy (anonymous):

Men / Women = 3/5 Women = Men * 5 / 3 Calculate Women. Then add Men to get total.